A current of 5.0A shows through an automobile headlight. How many coulombs of charge show through the headlamp in 2.00hrs
1
Expert's answer
2019-07-24T15:51:40-0400

Charge is given by

"Q=It"

where "I" is the current and "t" is the time.

Given "I=5.0\ A\ and \ t=2.00 hrs = 7200\ s"


"\therefore Q=2\times 7200\ C = 1.44\times 10^4\ C"

Thus the charge flowing through the headlamp is "1.44\times 10^4\ C"